# A Survey on SAR ship classification using Deep Learning

### v1

@@ V-C2 Local Radar Cross-Section
Local Radar Cross-Section (RCS) features refer to characteristics extracted from an image that are related to the radar scattering properties of specific image regions. These features are also combined with additional DL features in the literature to improve the performance of DL models [37].
@@ V-C3 Geometric Features
Geometric features can be extracted to utilize information such as length and width. During training with annotated data sets, these features can be drawn from the AIS information [32]. During operational classification, some bounding box algorithm can be used [41], as shown in Fig. 4, provided that the pixel per meter ratio is known. Length and width can be used to extract additional features [37], such as area, perimeter, aspect ratio, shape complex etc. Adding this information about geometric features can help DL models to better predict the ship classes.
@@ V-D Training Techniques
Although handcrafted features can be helpful, they require high-resolution images for effective extraction, which is often not the case, as seen in the OpenSARShip dataset. To address this challenge, researchers have explored novel deep learning techniques, further explained below.
@@ V-D1 Feature Fusion
A DL model trains on the features that it extracts, but when it comes to SAR ship classification, the challenge of data scarcity becomes a hurdle on the way to better results. To help the model understand the data well, several researchers propose to add hand-crafted features to the DL model [31, 43, 67, 60, 68]. Deep and handcrafted features are merged and the model classifies on this merged information, thus obtaining a better classification performance [37]. A merging strategy is schematized in Fig. 5.
@@ V-D2 Polarization fusion
SAR images can be acquired in four transmit-receive polarizations: VV, VH, HH, and HV. While most studies in the literature have focused solely on VV or VH polarizations, it is important to note that each polarization provides distinct information about the target. Utilizing multiple polarizations (e.g., VV-VH pairs) has been shown to improve classification performance. This is because combining information from different polarizations allows the model to learn more discriminative features, leading to enhanced classification accuracy. A common approach in the literature involves creating image pairs (e.g., VV-VH) and training two separate deep learning networks simultaneously, one for each polarization [51, 40, 46, 55, 28]. As illustrated in Fig. 6, the features extracted from these networks are then merged for final classification.
@@ V-D3 Metric Learning
DL models rely on feature extraction, which involves iteratively adjusting the model’s weights to optimize the extracted features. Ideally, for effective classification, the features corresponding to each class should cluster tightly in a low-dimensional space, while remaining well-separated from features of other classes. However, standard training procedures may not inherently prioritize this inter-class separation. In such cases, metric learning techniques are employed to explicitly address this challenge [30, 58, 34]. Metric learning aims to embed the extracted features into a low-dimensional space while maximizing the similarity between features belonging to the same class and minimizing the similarity between features from different classes. This is often achieved by utilizing distance metrics like cosine similarity or Euclidean distance (as illustrated in Fig. 7). Metric learning helps the model clearly distinguish between features of different classes, leading to superior performance. However, since the data is often naturally imbalanced, metric learning may be prone to overfitting. Therefore, it is crucial to balance the data beforehand or account for the imbalance when applying metric learning.
@@ V-D4 Multi-Scale/Pyramid Training
A DL model consists of several layers; each layer learns different scaled features, and the model makes the prediction based on the features of the last convolution layers [62, 61, 36, 33, 64, 40]. In pyramid training, see Fig. 8, the features from all the layers are directly used in decision-making by merging them using concatenation or other statistical measures (e.g. average, addition etc.).
@@ V-D5 Transfer-Learning and Domain Adaptation
SAR data are expensive and difficult to access, while DL models require large amounts of data. A deep neural network learns how to extract features and then applies classification to these learned features. Transfer learning can be beneficial in such scenarios. It allows leveraging other datasets to help a network learn feature extraction and then fine-tune the network with the SAR data. Transfer learning can be applied using various ways, see Fig. 9:
@@
    Merging optical and SAR imagery to see whether the model performs better at feature extraction, thus leading to improved overall performance [58, 45, 65, 63].
@@ V-D6 Learning Rate management
One crucial component of DL is the learning rate, which helps the network to update its weights efficiently. Traditionally, the learning rate is kept fixed during the entire training. However, as a deep neural network progresses through training iterations, it gradually learns information from the data. A constant learning rate in this scenario can slow down training and potentially cause the model to miss the point of lowest error. To address this issue, a strategy was introduced by [47] that changes the learning rate based on the number of iterations. This approach reduces the learning rate as training progresses, facilitating convergence towards the optimal weights. This technique not only reduces the training time but can also improve the overall performance of deep neural networks. In tasks like SAR ship classification, where data availability is limited, such dynamic learning rate adjustments can be particularly beneficial for achieving better performances.
@@ VI-A Model Architectures
    InceptionNet, also known as GoogLeNet [83], introduces the inception module, which allows for multi-scale processing by applying parallel convolutional filters of different sizes ($1\times 11 × 1$, $3\times 33 × 3$ and $5\times 55 × 5$) within the same layer. This architecture comprises 22 layers, with inception modules stacked sequentially to capture diverse spatial features. This design efficiently manages computational resources while enhancing the network’s ability to learn complex and varied patterns in the input data.
These models are often adapted for the SAR image data through transfer learning approaches. Additionally, generative adversarial networks have also been explored for potential applications in SAR ship classification [45, 63].
@@ VI-C1 Epochs
In deep learning, an epoch is a fundamental unit of training, defined as one complete iteration over the entire training set. During each epoch, the neural network processes every sample in the dataset, enabling the model to learn and adjust its parameters—such as weights and biases—based on the error between its predictions and the actual target values. The role of epochs is critical in the iterative optimization process, as each epoch provides an opportunity for the model to refine its internal representations and progressively minimize the loss function.
The importance of epochs lies in balancing the trade-off between underfitting and overfitting; a sufficient number of epochs ensures that the model captures the underlying patterns in the data, while avoiding excessive training that might lead to overfitting, where the model performs well on the training data but poorly on new, unseen data. Properly tuning the number of epochs, often through techniques like early stopping, is essential for achieving optimal model performance and generalization. The literature suggests a common range of 50–200 epochs for DL model convergence, but the optimal number ultimately depends on the chosen learning strategy [66, 77, 58, 63, 62, 61, 60].
@@ VI-C2 Batch Size
In deep learning, batch size refers to the number of training samples processed before updating the model’s parameters. It affects training stability and efficiency: smaller batch sizes offer more frequent updates and can improve generalization, but may introduce noise, while larger batch sizes provide more stable and accurate gradient estimates but require more memory and can slow down the training per epoch. Thus, selecting an appropriate batch size is crucial for balancing training speed, computational resources, and model performance. In the literature, batch sizes of 64, 32 and 16 are commonly used [67, 77, 64, 62, 61, 60, 49].
@@ VI-C3 Image Size
Selection of Image size plays a crucial role in determining the network computational requirements, memory usage, and ability to capture detailed features. Larger image sizes can provide more detailed spatial information, enhancing the model’s ability to detect fine-grained patterns, but they also significantly increase computational load and memory consumption. Conversely, smaller image sizes reduce computational demands and memory usage but may result in the loss of important visual details, potentially degrading the model performance. The literature commonly utilizes image sizes of $224\times 224224 × 224$, $64\times 6464 × 64$ and $128\times 128128 × 128$ [67, 66, 64, 63].
@@ VI-C4 Learning Rate
The learning Rate, see also Section V-D6 above, determines the step size at each iteration while moving towards a minimum of a loss function. It controls the weights of a DL network with respect to the loss gradient. A high learning rate can accelerate training by making large updates to the parameters, but risks overshooting the optimal values, leading to divergence or instability. Conversely, a low learning rate ensures more precise adjustments, promoting stable convergence, but can make the training process exceedingly slow and may get stuck in local minima. Therefore, selecting an appropriate learning rate is critical for achieving efficient and effective training. In the literature, the common learning rate is usually set in the range 0.0001 to 0.001 [67, 66, 58, 64, 63, 60]; moreover, learning rate management can be employed to train the model faster and more precisely [62, 61].
@@ VI-D Number of Classes
Due to the imbalanced dataset problem in SAR ship data, an increase in the number of classes results in fewer samples per class. For instance, in both the public datasets considered here (FUSARShip and OpenSARShip), over 80% of the data are concentrated in just five classes on average, with the remaining data sparsely distributed among other classes. This imbalance causes classes with fewer samples to have negligible impact during training, leading the model to overfit towards the more prevalent classes. Consequently, most research focuses on 3-5 classes; although some studies have explored using more than five classes, model performance significantly deteriorates as the number of classes increases.
@@ VII Outlook
This survey provides a comprehensive review of various techniques employed to train DL models for ship classification using SAR imagery. Through a systematic literature review, we meticulously assessed the content of 161 papers, rigorously narrowing down our selection to the most pertinent 49 studies. Their references are collected in Table II and their features are summarized in Section IV, offering an overview of their key elements and contributions to the field.
The literature reveals a broad range of DL techniques employed for SAR ship classification, as illustrated in Figure 2. These techniques include two types of architectures: CNNs, the earliest proposals, and state-of-the-art architectures such as ResNet, VGG, DenseNet etc., each tailored to leverage the unique attributes of SAR images.
@@
The FUSARShip and OpenSARShip datasets, along with other distinct data sources, are extensively used in the literature. These datasets are crucial for training and evaluating DL models. However, the limited availability of SAR images poses a significant challenge. To address this, data augmentation techniques have been proposed and are widely adopted. These techniques artificially increase the diversity and quantity of training data, which is essential for improving the performance and robustness of DL models. Fine-tuning pre-trained DL models on SAR-specific datasets has emerged as a successful strategy. This approach leverages the knowledge from pre-trained models and adapts it to the specific characteristics of SAR data, resulting in improved classification accuracy.
The issue of data imbalance has been addressed in the literature through two main solutions: adjusting train-test ratios and employing tailored loss functions. Additionally, some researchers propose the use of architecture search techniques like NAS for selecting the optimal DL architecture. This survey also highlights the significance of various training techniques, including feature fusion, polarization fusion, metric learning, multi-scale/pyramid learning, and learning rate management.
The findings of our survey underscore the importance of data quality and the choice of techniques in SAR ship classification. The integration of handcrafted features, effective data augmentation, and fine-tuning of pre-trained models are all critical factors contributing to the success of DL models in this domain. Building on the findings of this survey and acknowledging its limitations, several exciting avenues exist for future research in SAR ship classification using DL.

### v2

@@ V-C2 Local Radar Cross-Section
Local Radar Cross-Section (RCS) features describe scattering intensities within specific image regions. These features directly capture radar-domain priors that are not always learned effectively by DL models. When fused with learned features, local RCS provides physical interpretability and helps reduce confusion between ships with similar shapes but different scattering patterns [37]. Nonetheless, RCS features are sensitive to resolution and viewing geometry, which may limit generalization across datasets.
@@ V-C3 Geometric Features
Geometric features (such as length, width or related combinations) can be proficiently exploited for classification. During training with annotated datasets, these features can be drawn from the AIS information [32]. During operational classification, image processing algorithms can be applied to segment ship targets (e.g. through bounding box fit) and, provided that the resolution is known, estimate the geometric variables of interest [41], as shown in Fig. 8. Length and width can be used to extract additional features [37], such as area, perimeter, aspect ratio, shape etc. Adding this information can help DL models to improve the ship classes prediction.
@@
Overall, handcrafted features act as complementary signals rather than stand-alone solutions. HOG captures edge structure but is noise-sensitive, RCS injects radar-specific priors but depends on resolution, and geometric descriptors are highly discriminative yet prone to segmentation errors. Their effectiveness lies in feature fusion with DL models, especially under data scarcity and imbalance, where handcrafted cues can provide stability and interpretability.
@@ V-D Training Techniques
Although handcrafted features can be helpful, they require high-resolution images for effective extraction, which is often not the case, as seen in the OpenSARShip dataset. To address this challenge, researchers have explored novel DL techniques, further explained below.
@@ V-D1 Feature Fusion
a DL model trains on the features that it extracts, but when it comes to SAR ship classification, the challenge of data scarcity becomes a hurdle on the way to better results. To help the model understand the data well, several researchers propose to add handcrafted features to the DL model [31, 46, 70, 63, 71]. Deep and handcrafted features are merged and the model classifies on this merged information, thus obtaining a better classification performance [37]. A merging strategy is schematized in Fig. 9.
@@ V-D2 Polarization fusion
SAR images can be acquired in four transmit-receive polarizations: VV, VH, HH, and HV. While most studies in the literature have focused solely on VV or VH polarizations, it is important to note that each polarization provides distinct information about the target. Utilizing multiple polarizations has been shown to improve classification performance. This is because combining information from different polarizations allows the model to learn more discriminative features, leading to enhanced classification accuracy. A common approach in the literature involves creating image pairs and training two separate DL networks simultaneously, one for each polarization [54, 40, 49, 58, 28]. As illustrated in Fig. 10, the features extracted from these networks are then merged for final classification.
@@ V-D3 Metric Learning
DL models rely on feature extraction, which involves iteratively adjusting the model’s weights to optimize the extracted features. Ideally, for effective classification, the features corresponding to each class should cluster tightly in a low-dimensional space, while remaining well-separated from features of other classes. However, standard training procedures may not inherently prioritize this inter-class separation. In such cases, metric learning techniques are employed to explicitly address this challenge [30, 61, 34]. Metric learning aims to embed the extracted features into a low-dimensional space while maximizing the similarity between features belonging to the same class and minimizing the similarity between features from different classes. This is often achieved by utilizing distance metrics like cosine similarity or Euclidean distance, as illustrated in Fig. 11. Metric learning helps the model clearly distinguish between features of different classes, leading to superior performance. However, since the data is often naturally imbalanced, metric learning may be prone to overfitting. Therefore, it is crucial to balance the data beforehand or account for the imbalance when applying metric learning.
@@ V-D4 Multi-Scale/Pyramid Training
A DL model consists of several layers; each layer learns different scaled features, and the model makes the prediction based on the features of the last convolution layers [65, 64, 36, 33, 67, 40, 88]. In pyramid training, see Fig. 12, the features from all the layers are directly used in decision-making by merging them using concatenation or other statistical measures (e.g. average, addition etc.).
@@ V-D5 Transfer-Learning and Domain Adaptation
SAR data are expensive and difficult to access, while DL models require large amounts of data. A deep neural network learns how to extract features and then applies classification to these learned features. Transfer learning can be beneficial in such scenarios. It allows leveraging other datasets to help a network learn feature extraction and then fine-tune the network with the SAR data. Transfer learning can be applied in various ways, see Fig. 13:
@@
    Merging optical and SAR imagery to see whether the model performs better at feature extraction, thus leading to improved overall performance [61, 48, 68, 66, 86].
@@ V-D6 Learning Rate management
One crucial component of DL is the learning rate, which helps the network to update its weights efficiently. Traditionally, the learning rate is kept fixed during the entire training. However, as a deep neural network progresses through training iterations, it gradually learns information from the data. A constant learning rate in this scenario can slow down training or potentially cause the model to miss the point of lowest error. To address this issue, a strategy was introduced in [50] that changes the learning rate based on the number of iterations. This approach reduces the learning rate as training progresses, facilitating convergence towards the optimal weights. This technique not only reduces the training time but can also improve the overall performance of deep neural networks. In tasks like SAR ship classification, where data availability is limited, such dynamic learning rate adjustments can be particularly beneficial for achieving better performances.
@@ VI-A Model Architectures
    InceptionNet, also known as GoogLeNet [108], introduces the inception module, which allows for multi-scale processing by applying parallel convolutional filters of different sizes ($1\times 1$, $3\times 3$ and $5\times 5$) within the same layer. This architecture comprises 22 layers, with inception modules stacked sequentially to capture diverse spatial features. This design efficiently manages computational resources while enhancing the network’s ability to learn complex and varied patterns in the input data.
These models are often adapted for SAR data processing through transfer learning approaches. Additionally, generative adversarial networks have also been explored for potential applications in SAR ship classification [48, 66].
@@ VI-C1 Epochs
In DL, an epoch is a fundamental unit of training, defined as one complete iteration over the entire training set. During each epoch, the neural network processes every sample in the dataset, enabling the model to learn and adjust its parameters—such as weights and biases—based on the error between its predictions and the actual target values. The role of epochs is critical in the iterative optimization process, as each epoch provides an opportunity for the model to refine its internal representations and progressively minimize the loss function.
The importance of epochs lies in balancing the trade-off between underfitting and overfitting; a sufficient number of epochs ensures that the model captures the underlying patterns in the data, while avoiding excessive training that might lead to overfitting, where the model performs well on the training data but poorly on new, unseen data. Properly tuning the number of epochs, often through techniques such as early stopping, is essential to achieve optimal model performance and generalization. The literature suggests a common range of 50–200 epochs for DL model convergence, but the optimal number ultimately depends on the chosen learning strategy [69, 76, 61, 66, 65, 64, 63].
@@ VI-C2 Batch Size
In DL, batch size refers to the number of training samples processed before updating the model parameters. It affects training stability and efficiency: smaller batch sizes offer more frequent updates and can improve generalization, but may introduce noise, while larger batch sizes provide more stable and accurate gradient estimates but require more memory and can slow down the training per epoch. Thus, selecting an appropriate batch size is crucial for balancing training speed, computational resources, and model performance. In the literature, batch sizes of 64, 32 and 16 are commonly used [70, 76, 67, 65, 64, 63, 52].
@@ VI-C3 Image Size
Selecting image size plays a crucial role in determining the network computational requirements, memory usage, and ability to capture detailed features. Larger image sizes can provide more detailed spatial information, enhancing the model’s ability to detect fine-grained patterns, but they also significantly increase computational load and memory consumption. Conversely, smaller image sizes reduce computational demands and memory usage but may result in the loss of important visual details, potentially degrading the model performance. The literature commonly utilizes image sizes of $224\times 224$, $64\times 64$ and $128\times 128$ [70, 69, 67, 66].
@@ VI-C4 Learning Rate
The Learning Rate (see also Section V-D6) determines the step size at each iteration while moving towards a minimum of a loss function. It controls the weights of a DL network with respect to the loss gradient. A high learning rate can accelerate training by making large updates to the parameters, but risks overshooting the optimal values, leading to divergence or instability. Conversely, a low learning rate ensures more precise adjustments, promoting stable convergence, but can make the training process exceedingly slow and may get stuck in local minima. Therefore, selecting an appropriate learning rate is critical for achieving efficient and effective training. In the literature, the common learning rate is usually set in the range 0.0001 to 0.001 [70, 69, 61, 67, 66, 63]; moreover, learning rate management can be employed to train the model faster and more precisely [65, 64].
@@ VI-D Number of Classes
Due to the imbalanced dataset problem in SAR ship data, an increase in the number of classes results in fewer samples per class. For instance, in both the public datasets considered here (FUSARShip and OpenSARShip), over 80% of the data are concentrated in just five classes on average, with the remaining data sparsely distributed among other classes. This imbalance causes classes with fewer samples to have negligible impact during training, leading the model to overfit towards the more prevalent classes. Consequently, most research focuses on 3-5 classes (see tables-V, VI); although a few studies have explored the usage of more than five classes, it is common experience that the model performance significantly deteriorates as the number of classes increases.
TABLE V: SAR ship-classification results on public datasets with $\leq\!3$ target classes. Yellow cells mark the best Overall Accuracy (OA) and F1 within each class-count block. Method abbreviations: CNN—Convolutional Neural Network; DA—Domain Adaptation; FF—Feature Fusion; GAN—Generative Adversarial Network; HC—Hand-Crafted features; KD—Knowledge Distillation; ML—Metric Learning; MS—Multi-Scale; NAS—Neural Architecture Search; PF—Polarization Fusion; Dataset abbreviations: CSK—COSMO-SkyMed; FS—FUSAR-Ship; GF-3—Gaofen-3; OS—OpenSARShip; TX—TerraSAR-X.
Method Dataset #Cls. OA [%] F1 [%] PF (2021)[41] GF-3 2 96.55 93.88 HC FF (2023)[94] OS 2 88.87 83.49 CNN (2018)[22] CSK 3 97.66 97.80 DA+TL (2019)[22] TX 3 98.52 97.15 Inc. Angle (2019)[25] CSK 3 98.13 Upsampling (2019)[26] OS 3 81.56 ML (2020)[30] OS 3 83.67 GAN (2020)[48] GF-3 3 95.00 FF (2021)[31] OS 3 77.69 AIS+CNN (2021)[32] OS 3 93.30 93.73 ML (2021)[34] OS 3 88.97 Attn. FF (2021)[36] OS 3 92.66 FF (2021)[37] OS 3 78.00 FF (2021)[37] FS 3 86.86 MS+PF (2021)[40] OS 3 79.25 77.62 FF (2021)[44] OS 3 77.69 MS+HC FF (2022)[46] OS 3 78.15 75.04 FF (2022)[47] TX 3 88.89 FF (2022)[47] FS 3 89.11 PF (2022)[49] OS 3 88.00 PF (2022)[52] OS 3 88.80 86.81 NAS (2022)[53] OS 3 88.02 PF+HC (2022)[54] OS 3 79.84 78.00 DA (2022)[55] TX 3 87.67 NAS (2023)[56] OS 3 82.06 PF (2023)[58] OS 3 81.28 MS+PF (2023)[59] OS 3 88.93 86.54 Novel CNN (2023)[65] GF-3 3 89.79 89.71 Attn. Dense GAN (2023)[66] FS 3 92.00 MS (2023)[67] OS 3 82.23 80.91 Dual-Pol (2024)[69] OS 3 87.62 86.74 DA (2024)[73] FS 3 75.94 Novel Arch. (2024)[75] OS 3 81.86 80.75 Novel Arch. (2024)[75] FS 3 75.18 73.05 PF (2024)[76] OS 3 87.13 86.27 PF (2024)[77] OS 3 89.40 86.89 FF (2024)[78] OS 3 82.00 82.00 FF (2024)[78] FS 3 90.00 89.00 MS (2024)[79] OS 3 85.66 84.90 MS (2024)[79] FS 3 85.26 86.17 B-CNN (2024)[80] OS 3 76.30 76.30 FM (2024)[81] OS 3 79.02 KD (2024)[82] OS 3 80.03 Imbal. (2024)[83] FS+OS 3 74.60 67.50 CV-DL (2025)[85] GF-3 3 72.22 DA (2025)[86] FS 3 95.33 DA (2025)[86] TX 3 89.33 MS FF (2025)[88] OS 3 80.09 79.27 Super-Res (2025)[89] OS 3 72.94 61.53 Super-Res (2025)[90] OS 3 87.68
TABLE VI: SAR ship-classification results on public datasets with >3>\!3 target classes. Abbreviations as in Table V.
Method Dataset #Cls. OA. [%] F1 [%] CNN + Spatial Attention (2018)[23] OS 4 84.00 CNN (2019)[27] P2 4 65.70 Deep FF (2019)[28] OS 4 87.04 FF (2021)[39] OS 4 82.35 82.19 DA (2022)[55] FS 4 82.33 NAS (2023)[56] FS 4 63.90 HC (2024)[70] OS 4 87.50 86.00 LSTM + CNN (2024)[74] FS 4 65.58 Complex-Valued DL (2025)[85] GF-3 4 69.01 CNN (2018)[21] TX 5 94.00 Metric Learning (2021)[34] OS 5 68.16 PF (2022)[52] OS 5 66.90 57.54 Novel CNN (2023)[57] FS 5 98.71 97.69 MS + PF (2023)[59] OS 5 67.03 57.89 Data Augmentation (2024)[72] FS 5 72.19 71.97 PF (2024)[77] OS 5 67.83 58.01 DA (2024)[68] FS 5 91.22 MS (2024)[79] FS 5 84.24 84.36 Metric Learning (2024)[61] FS 5 78.57 MS + PF (2021)[40] OS 6 56.66 51.38 Sampling (2021)[43] OS 6 80.58 NAS (2022)[54] OS 6 56.73 PF + HC (2022)[54] OS 6 56.83 51.82 PF (2023)[58] OS 6 58.68 MS (2023)[60] OS 6 81.83 MS (2023)[67] OS 6 60.76 59.89 Dual-Pol (2024)[69] OS 6 64.26 58.44 PF (2024)[76] OS 6 65.97 58.63 MS FF (2025)[88] OS 6 60.63 58.29 Super-Res (2025)[90] OS 6 69.71 Super-Res (2025)[92] OS 6 65.40 FF (2021)[37] FS 7 86.86 MS + HC FF (2022)[46] FS 7 86.69 86.58 HC FF (2023)[63] FS 7 87.23 86.69 MS (2023)[67] FS 7 90.08 90.12 Feature Manip. (2024)[81] FS 7 87.59 Knowledge Distill. (2024)[82] FS 7 86.07 HC (2024)[71] FS 7 87.94 Deep FF (2025)[84] FS 8 76.12 56.67 Curriculum Learning (2025)[87] FS 9 68.34 55.49 Over-Sampling (2025)[93] FS 9 52.87 Side-Lobe Elimin. (2021)[38] FS 10 98.40 97.96 Siamese Net (2021)[51] OS 16 97.97
TABLE VII: Challenges in SAR ship classification and practical remedies.
Challenge Remedies / notes
Data scarcity High-quality open SAR curation; AIS–SAR mapping for labels; strong augmentations; transfer learning.
Class imbalance Rebalanced splits; focal/Class Balanced loss; mixup/cutmix/feature-space-oversampling; report F1 alongside OA, curriculum learning.
Domain shift (sensor/scene) Normalization; domain adaptation; test-time adaptation; confidence calibration.
Small targets / clutter Multi-scale features; higher input resolution; denoising; polarization/feature fusion, super resolution.
Reproducibility Report splits/seeds/hyperparams; share CSV of results.
@@ VII Outlook
This survey provides a comprehensive review of various techniques employed to train DL models for ship classification using SAR imagery. Through a systematic literature review, we assessed the content of 187 papers, rigorously narrowing down our selection to the most pertinent 74 studies. Their references are collected in Table II and their features are summarized in Section IV, offering an overview of their key elements and contributions to the field.
The literature reveals a broad range of DL techniques employed for SAR ship classification, as illustrated in Figure 2. These techniques include CNNs, in the earliest proposals, and state-of-the-art architectures such as ResNet, VGG, DenseNet etc., each one tailored to leverage the unique attributes of SAR images.
@@
The FUSARShip and OpenSARShip datasets, along with other distinct data sources, are extensively used in the literature. These datasets are crucial for training and evaluating DL models. However, the limited availability of SAR images poses a significant challenge. To address this, data augmentation techniques have been proposed and widely adopted. These techniques artificially increase the diversity and quantity of training data, which is essential for improving the performance and robustness of DL models. Fine-tuning pre-trained DL models on SAR-specific datasets has emerged as a successful strategy. This approach leverages the knowledge from pre-trained models and adapts it to the specific characteristics of SAR data, resulting in improved classification accuracy.
The issue of data imbalance has been addressed in the literature through two main solutions: adjusting train-test datasets ratios and employing tailored loss functions. Additionally, some researchers propose the use of architecture search techniques like NAS for selecting the optimal DL architecture. This survey also highlights the significance of various training techniques, including feature fusion, polarization fusion, metric learning, multi-scale/pyramid learning, and learning rate management.
The findings of our survey underscore the importance of data quality and the choice of techniques in SAR ship classification. The integration of handcrafted features, effective data augmentation, and fine-tuning of pre-trained models are all critical factors contributing to the success of DL models in this domain. Building on the findings of this survey and acknowledging the limitations of current research, several exciting avenues for future work in SAR ship classification using DL are identified and described in the following subsections.